# Enhancing Reasoning for Diffusion LLMs via Distribution Matching Policy Optimization

### v1

@@ 3 Distribution Matching Policy Optimization
3.1 From Reward Maximization to Distribution Matching
Figure 2: Illustration of relative entropy (mode-seeking) and cross-entropy (mass-covering) for fitting a target $p_{*}$ ($\mathcal{G}$ is the set of Gaussian distributions)
@@ 3.1 From Reward Maximization to Distribution Matching
$\max_{\theta}\operatorname{\mathbb{E}}_{{\bm{q}}\sim\mathcal{D}}\left[\operatorname{\mathbb{E}}_{p_{\theta}({\bm{o}}|{\bm{q}})}[r({\bm{q}},{\bm{o}})]-\alpha\operatorname{KL}(p_{\theta}(\cdot|{\bm{q}})\|p_{\mathrm{ref}}(\cdot|{\bm{q}}))\right].$ (4)
However, existing techniques over-focus on finding and optimizing the reward mode and adopt many heuristic techniques to accelerate the mode searching process, neglecting the exploration of the entire distribution landscape, and often result in model mode collapse or reward hacking, causing the model to produce undesirable responses (Weng, 2024). A simple fix to this issue and to encourage diverse model responses is to enforce the optimality of the target policy distribution during the training. It can be shown that the optimal sequence distribution that solves the problem ˜4 is the following reward-tilted distribution:
$p_{*}({\bm{o}}|{\bm{q}})=\frac{1}{Z({\bm{q}})}p_{\mathrm{ref}}({\bm{o}}|{\bm{q}})\mathrm{e}^{r({\bm{q}},{\bm{o}})/\alpha},\quad\text{where}~Z({\bm{q}})=\sum_{\bm{o}}p_{\mathrm{ref}}({\bm{o}}|{\bm{q}})\mathrm{e}^{r({\bm{q}},{\bm{o}})/\alpha}.$ (5)
That is to say, we want to use the optimal sequential distribution $p_{*}({\bm{o}}|{\bm{q}})$ as the supervision signal throughout the learning process, so that we can learn a dLLM policy ${\bm{\pi}}_{\theta}$ which produces a sequence distribution $p_{\theta}$ matching $p_{*}$. We can thus obtain a policy that not only explores the dominant reward mode, but is guaranteed to sample other high-reward trajectories with a likelihood proportional to the reward value. This motivates us to consider the following task of policy distribution matching,
Policy Distribution Matching Learning: Given a pretrained dLLM policy $\bm{\pi}_{\mathrm{ref}}({\bm{o}}|{\bm{q}})$ that samples from a distribution $p_{\mathrm{ref}}({\bm{o}}|{\bm{q}})$, a reward function $r:({\bm{q}},{\bm{o}})\mapsto\mathbb{R}$, a set of prompts $\mathcal{D}$, and temperature $\alpha>0$, learn a dLLM policy ${\bm{\pi}}_{\theta}({\bm{o}}|{\bm{q}})$ to produce the desired optimal sequence distribution $p_{*}({\bm{o}}|{\bm{q}})$ in ˜5 by optimizing the following objective: $\min_{{\bm{\pi}}_{\theta}}\operatorname{\mathbb{E}}_{{\bm{q}}\sim\mathcal{D}}\mathcal{F}(p_{\theta}(\cdot|{\bm{q}}),p_{*}(\cdot|{\bm{q}})).$ (6)
@@
To address this issue, we consider a series of new objectives $\mathcal{F}$ with more desirable convergence guarantees that steadily lead to optimization towards the desired sequence distribution, and propose Distribution Matching Policy Optimization (DMPO) (Alg.˜1), which targets matching the entire reward-tilted policy distribution. In Sec.˜3.2, we introduce weighted denoising cross-entropy, a scalable implementation of the forward KL using importance sampling. In Sec.˜3.3 and Sec.˜3.4, we discuss an important failure case of forward KL with small training batch size, and propose a series of novel techniques such as weight baseline subtraction (Sec.˜3.3) and weighted direct discriminative optimization (Sec.˜3.4) to address it.
Algorithm 1 Distribution Matching Policy Optimization (DMPO)
1:Training dataset $\mathcal{D}$, number of prompts per batch $B$, number of rollouts per prompt $N$, frequency for sampling buffer $F$, model policy ${\bm{\pi}}_{\theta}$.
2:for $\mathtt{step}=0,1,2,$... do
3:  if $\mathtt{step}~\mathrm{mod}~F=0$ then $\triangleright$ Prepare the buffer using the current policy, denoted ${\bm{\pi}}_{v}$.
4:   Sample $B$ prompts $\{{\bm{q}}^{(i)}\}_{1\leq i\leq B}$ from the dataset $\mathcal{D}$.
5:   for $1\leq i\leq B$ (in parallel, with gradient computation disabled) do
6:     Sample NN orders and generate NN rollouts $\{{\bm{o}}^{(i,n)}\}_{1\leq n\leq N}$ conditional on prompt ${\bm{q}}^{(i)}$.
7:     Evaluate reward and compute weights $w({\bm{o}}^{(i,n)}|{\bm{q}}^{(i)};{\bm{\sigma}}^{(i,n)})$ according to ˜10.
8:     Compute the weight baseline according to ˜13, ˜14, or ˜15, and obtain the real weights $w_{(}{\bm{o}}^{(i,n)}|{\bm{q}}^{(i)};{\bm{\sigma}}^{(i,n)})$ according to ˜12.
9:  For each ${\bm{o}}^{(i,n)}$, sample a mask assignment and obtain ${\widetilde{\bm{o}}}^{(i,n)}$.
10:  Feed all pairs of $({\bm{q}}^{(i)},{\widetilde{\bm{o}}}^{(i,n)})$ into ${\bm{\pi}}_{\theta}$ and compute the WDCE loss ˜11, then update $\theta$. return ${\bm{\pi}}_{\theta}$
@@ 3.2 Weighted Denoising Cross-entropy
Unlike the reverse KL objective considered by many existing works, which are known to be prone to mode seeking and collapse, one alternative choice is to use the forward KL divergence (or cross-entropy, CE) for the functional, i.e., $\mathcal{F}(p_{\theta},p_{*})=\operatorname{KL}(p_{*}\|p_{\theta})$, which tends to cover all the modes of the optimal distribution and can retain the response diversity. The CE loss is also widely used in another domain known as stochastic optimal control (SOC) (Domingo-Enrich et al., 2024; 2025), which is also closely connected with our work. This amounts to solving the following task,
$\min_{\theta}\operatorname{\mathbb{E}}_{{\bm{q}}\sim\mathcal{D}}\operatorname{\mathbb{E}}_{p_{*}({\bm{o}}|{\bm{q}})}\left[\log\dfrac{p_{*}({\bm{o}}|{\bm{q}})}{p_{\theta}({\bm{o}}|{\bm{q}})}\right].$ (7)
However, objective ˜7 is not directly amenable to practical implementation as we do not have access to real samples from the $p_{*}$, nor can we exactly compute $\log p_{*}$ due to the presence of the unknown partition function $Z({\bm{q}})$. To bypass this issue, we draw inspiration from the recent work masked diffusion neural sampler (MDNS, Zhu et al. (2025g)), which proposes a training framework for learning a masked diffusion neural sampler with stochastic optimal control and cross-entropy minimization. While targeting a different task, the core of MDNS resides in solving the same distribution matching problem with cross-entropy loss, and it proposes a practically implementable and scalable variant of ˜7, named weighted denoising cross-entropy (WDCE) loss. The central idea is to introduce a reference policy and leverage importance sampling, so that we can treat i.i.d. samples as importance-weighted samples from $p_{*}$. Taking advantage of this approach, we now derive WDCE for the purpose of dLLM policy learning.
First, given the relationship between the policy output and sequence distribution of the masked dLLM ˜1, it is clear that we can match the correct target sequence distribution $p_{*}({\bm{o}}|{\bm{q}})$ as long as we train $p_{\theta}({\bm{o}}|{\bm{q}};{\bm{\sigma}})$ to match the order-specific ones, i.e., $p_{*}({\bm{o}}|{\bm{q}};{\bm{\sigma}})$, given by
@@
Importance weight $w({\bm{o}}|{\bm{q}};{\bm{\sigma}})$. We simplify it with the pretrained model and the reward:
$w({\bm{o}}|{\bm{q}};{\bm{\sigma}})=\frac{1}{Z({\bm{q}})}\frac{p_{\mathrm{ref}}({\bm{o}}|{\bm{q}};{\bm{\sigma}})}{p_{v}({\bm{o}}|{\bm{q}};{\bm{\sigma}})}\mathrm{e}^{\frac{r({\bm{q}},{\bm{o}})}{\alpha}}\propto\exp\left(\frac{r({\bm{q}},{\bm{o}})}{\alpha}+\log\frac{p_{\mathrm{ref}}({\bm{o}}|{\bm{q}};{\bm{\sigma}})}{p_{v}({\bm{o}}|{\bm{q}};{\bm{\sigma}})}\right)=:\mathrm{e}^{\ell({\bm{o}}|{\bm{q}};{\bm{\sigma}})}.$ (10)
Recall that the order-specific probability of a sequence is computed via ˜1. To ensure that the sample distribution after importance sampling is valid and normalized, we keep track of the log weights $\ell({\bm{o}}|{\bm{q}};{\bm{\sigma}})$, and taking softmax among those corresponding to the same prompt $\bm{q}$} to compute the real weight $w({\bm{o}}|{\bm{q}};{\bm{\sigma}})$. This is equivalent to estimating the unknown partition function $Z({\bm{q}})$ using an empirical estimator of the following expectation:
$Z({\bm{q}})=\operatorname{\mathbb{E}}_{\bm{\sigma}}\operatorname{\mathbb{E}}_{p_{v}({\bm{o}}|{\bm{q}};{\bm{\sigma}})}\left[\frac{p_{\mathrm{ref}}({\bm{o}}|{\bm{q}};{\bm{\sigma}})}{p_{v}({\bm{o}}|{\bm{q}};{\bm{\sigma}})}\mathrm{e}^{r({\bm{q}},{\bm{o}})/\alpha}\right],$
@@
Sequence log probability $\log p_{v}({\bm{o}}|{\bm{q}})$. Unlike the case of LLM, the exact sequence log probability is intractable due to the presence of expectation over the random order $\bm{\sigma}$}. However, similar to the training of dLLM, we can leverage the negative ELBO ˜2 as a surrogate. Combined with the importance weight $w({\bm{o}}|{\bm{q}};{\bm{\sigma}})$, we introduce the weighted denoising cross-entropy (WDCE) loss for dLLM policy distribution matching:
$\min_{\theta}\underset{{\bm{q}}\sim\mathcal{D}}{\operatorname{\mathbb{E}}}\,\underset{{\bm{\sigma}}}{\operatorname{\mathbb{E}}}\,\underset{p_{v}({\bm{o}}|{\bm{q}};{\bm{\sigma}})}{\operatorname{\mathbb{E}}}\bigg\{w({\bm{o}}|{\bm{q}};{\bm{\sigma}})\underset{m\sim\operatorname{Unif}\{1,...,|{\bm{o}}|\}}{\operatorname{\mathbb{E}}}\bigg[\frac{|{\bm{o}}|}{m}\underset{\mu_{m}({\widetilde{\bm{o}}}|{\bm{o}})}{\operatorname{\mathbb{E}}}\sum_{d:\widetilde{o}_{d}=\mathsf{M}}-\log{\bm{\pi}}_{\theta}({\widetilde{\bm{o}}}|{\bm{q}})_{d,o_{d}}\bigg]\bigg\}.$ (11)
Notably, this loss highly resembles the DCE loss used in pre-training and the SFT phase of dLLM. One major difference is that instead of using i.i.d. samples from $p_{*}$, we use importance sampling to weight samples from $p_{v}$ and obtain a valid training objective with theoretical guarantees. WDCE differs significantly from other popular RL training techniques such as PPO/GRPO in two key aspects.
WDCE is an off-policy loss. The WDCE loss remains valid as the model parameter $\theta$ gets updated, since both the sampling policy $p_{v}$ and the important sampling target policy $p_{*}$ are independent of the current model policy $p_{\theta}$. This allows us to save generated rollouts in a replay buffer and reuse them for multiple training updates, without worrying excessively about numerical instability, leading to improved sample efficiency. On the other hand, for on-policy methods, to use a replay buffer, one would need to estimate importance weights with respect to the current model policy $p_{\theta}({\bm{o}}|{\bm{q}};{\bm{\sigma}})$, i.e., $\frac{p_{\theta}({\bm{o}}|{\bm{q}};{\bm{\sigma}})}{p_{v}({\bm{o}}|{\bm{q}};{\bm{\sigma}})}$. Different from the case of LLM, where such estimation can be done in one model forward pass, an accurate estimation in dLLM per training update is expensive, rendering the on-policy method less efficient. Moreover, for large models, where rollout generation and sequence likelihood estimation are typically handled by different implementations (such as vLLM and FSDP), this could lead to more nuanced, hard-to-detect biases that secretly undermine the algorithm’s performance (Yao et al., 2025). With WDCE, we are largely free of such concerns.
WDCE is a forward loss. Unlike the GRPO-style of algorithms that typically require keeping track of the entire rollout trajectories, WDCE leverages the forward noising process in training, which is a characteristic unique to diffusion LLMs. Once we obtain the final samples and their associated weights, we can discard the trajectories and perform training using the cheap forward process by randomly masking the data. This implies that the training speed when using WDCE largely depends on the model inference speed. With the advances of dLLM efficient inference techniques such as fast decoding algorithms and KV-cache techniques (Ma et al., 2025; Hu et al., 2025; Wu et al., 2025; Liu et al., 2025b), WDCE could also enjoy a great boost in efficiency. This method also effectively utilizes dLLM’s potential in surpassing LLMs in inference throughput, distinguishing it from other RL baselines that merely adapt LLM algorithms to dLLM. We defer a more detailed discussion of such properties to Sec.˜B.1.
3.3 Effective Training with Negative Gradient Insertion
While theoretically, minimizing the WDCE loss ˜11 probably leads to convergence of the model sequence distribution to $p_{*}({\bm{o}}|{\bm{q}})$, this could face practicality issues in real implementation due to the often limited number of rollouts generated per prompt. Ideally, we would want to promote the likelihood of “good” responses while decreasing those of “bad” responses. However, with WDCE, any response $\bm{o}$} will be associated with a positive weight $w({\bm{o}}|{\bm{q}};{\bm{\sigma}})$ due to the softmax operation, which may lead to ineffective learning in the low-batch-size scenario.

### v2

@@ 3 Distribution Matching Policy Optimization
3.1 Reward Maximization $\to$ Distribution Matching
Figure 2: Illustration of relative entropy (mode-seeking) and cross-entropy (mass-covering) for fitting a target $p_{*}$ ($\mathcal{G}$ is the set of Gaussian distributions).
@@ 3.1 Reward Maximization $\to$ Distribution Matching
$\max_{\theta}\operatorname{\mathbb{E}}_{{\bm{q}}\sim\mathcal{D}}\big[\operatorname{\mathbb{E}}_{p_{\theta}({\bm{o}}|{\bm{q}})}r({\bm{q}},{\bm{o}})-\alpha\operatorname{KL}(p_{\theta}(\cdot|{\bm{q}})\|p_{\mathrm{ref}}(\cdot|{\bm{q}}))\big].$ (4)
However, existing techniques over-focus on finding and optimizing the reward mode and adopt many heuristic techniques to accelerate the mode searching process, neglecting the exploration of the entire distribution landscape, and often result in model mode collapse or reward hacking, causing the model to produce undesirable responses (Weng, 2024). A simple fix to this issue and to encourage diverse model responses is to enforce the optimality of the target policy distribution during the training. It can be shown that the optimal sequence distribution that solves the problem 4 is the following reward-tilted distribution:
$p_{*}({\bm{o}}|{\bm{q}})=\frac{1}{Z({\bm{q}})}p_{\mathrm{ref}}({\bm{o}}|{\bm{q}})\mathrm{e}^{r({\bm{q}},{\bm{o}})/\alpha},$ (5)
where $Z({\bm{q}})=\sum_{\bm{o}}p_{\mathrm{ref}}({\bm{o}}|{\bm{q}})\mathrm{e}^{r({\bm{q}},{\bm{o}})/\alpha}$. That is to say, we want to use the optimal sequential distribution $p_{*}({\bm{o}}|{\bm{q}})$ as the supervision signal throughout the learning process, so that we can learn a dLLM policy ${\bm{\pi}}_{\theta}$ which produces a sequence distribution $p_{\theta}$ matching $p_{*}$. We can thus obtain a policy that not only explores the dominant reward mode, but is guaranteed to sample other high-reward trajectories with a likelihood proportional to the reward value. This motivates us to consider the following task:
Policy Distribution Matching Learning: Given a pretrained dLLM policy $\bm{\pi}_{\mathrm{ref}}({\bm{o}}|{\bm{q}})$ that samples from a distribution $p_{\mathrm{ref}}({\bm{o}}|{\bm{q}})$, a reward function $r:({\bm{q}},{\bm{o}})\mapsto\mathbb{R}$, a set of prompts $\mathcal{D}$, and temperature $\alpha>0$, learn a dLLM policy ${\bm{\pi}}_{\theta}({\bm{o}}|{\bm{q}})$ to produce the desired optimal distribution $p_{*}({\bm{o}}|{\bm{q}})$ 5 by optimizing the following objective: $\min_{{\bm{\pi}}_{\theta}}\operatorname{\mathbb{E}}_{{\bm{q}}\sim\mathcal{D}}\mathcal{F}(p_{\theta}(\cdot|{\bm{q}}),p_{*}(\cdot|{\bm{q}})).$ (6)
@@
To address this issue, we consider a series of new objectives $\mathcal{F}$ with more desirable convergence guarantees that steadily lead to optimization towards the desired sequence distribution, and propose Distribution Matching Policy Optimization (DMPO) (Alg. 1), which targets matching the entire reward-tilted policy distribution. In Sec. 3.2, we introduce weighted denoising cross-entropy (WDCE), a scalable implementation of the forward KL using importance sampling. In Secs. 3.3 and 3.4, we discuss an important failure case of forward KL with small training batch size, and propose a series of novel techniques such as weight baseline subtraction (Sec. 3.3) and weighted direct discriminative optimization (Sec. 3.4) to address it.
@@ 3.2 Weighted Denoising Cross-entropy
Unlike the reverse KL objective considered by many existing works, which are known to be prone to mode seeking and collapse, one alternative choice is to use the forward KL divergence (or cross-entropy, CE) for the functional, i.e., $\mathcal{F}(p_{\theta},p_{*})=\operatorname{KL}(p_{*}\|p_{\theta})$, which tends to cover all the modes of the optimal distribution and can retain the response diversity. The CE loss is also widely used in another domain known as stochastic optimal control (SOC) (Domingo-Enrich et al., 2024, 2025), which is also closely connected with our work. This amounts to solving the following task,
$\min_{\theta}\operatorname{\mathbb{E}}_{{\bm{q}}\sim\mathcal{D}}\operatorname{\mathbb{E}}_{p_{*}({\bm{o}}|{\bm{q}})}\Big[\log\dfrac{p_{*}({\bm{o}}|{\bm{q}})}{p_{\theta}({\bm{o}}|{\bm{q}})}\Big].$ (7)
However, objective 7 is not directly amenable to practical implementation, as we do not have access to real samples from the $p_{*}$, nor can we exactly compute $\log p_{*}$ due to the presence of the unknown partition function $Z({\bm{q}})$. To bypass this issue, we draw inspiration from the recent work masked diffusion neural sampler (MDNS, Zhu et al. (2025g)), which proposed a training framework for learning a neural sampler based on MDM with stochastic optimal control and cross-entropy minimization. While targeting a different task, the core of MDNS resides in solving the same distribution matching problem with cross-entropy loss, and it proposed a practically implementable and scalable variant of 7 named weighted denoising cross-entropy (WDCE) loss. The central idea is to introduce a reference policy and leverage importance sampling, so that we can treat i.i.d. samples as importance-weighted samples from $p_{*}$. Taking advantage of this approach, we now derive WDCE for the purpose of dLLM policy learning.
First, given the relationship between the policy output and sequence distribution of the masked dLLM 1, it is clear that we can match the correct target sequence distribution $p_{*}({\bm{o}}|{\bm{q}})$ as long as we train $p_{\theta}({\bm{o}}|{\bm{q}};{\bm{\sigma}})$ to match the order-specific ones, i.e., $p_{*}({\bm{o}}|{\bm{q}};{\bm{\sigma}})$, given by
@@
Importance weight $w({\bm{o}}|{\bm{q}};{\bm{\sigma}})$
We simplify it with the pretrained model and the reward:
$w({\bm{o}}|{\bm{q}};{\bm{\sigma}})=\frac{1}{Z({\bm{q}})}\frac{p_{\mathrm{ref}}({\bm{o}}|{\bm{q}};{\bm{\sigma}})}{p_{v}({\bm{o}}|{\bm{q}};{\bm{\sigma}})}\mathrm{e}^{\frac{r({\bm{q}},{\bm{o}})}{\alpha}}$
$\propto$ $\exp\Big(\frac{r({\bm{q}},{\bm{o}})}{\alpha}+\log\frac{p_{\mathrm{ref}}({\bm{o}}|{\bm{q}};{\bm{\sigma}})}{p_{v}({\bm{o}}|{\bm{q}};{\bm{\sigma}})}\Big)=:\mathrm{e}^{\ell({\bm{o}}|{\bm{q}};{\bm{\sigma}})}.$ (10)
Recall that the order-specific probability of a sequence is computed via 1. To ensure that the sample distribution after importance sampling is valid and normalized, we keep track of the log weights $\ell({\bm{o}}|{\bm{q}};{\bm{\sigma}})$, and taking softmax among those corresponding to the same prompt $\bm{q}$} to compute the real weight $w({\bm{o}}|{\bm{q}};{\bm{\sigma}})$. This is equivalent to estimating the unknown partition function $Z({\bm{q}})$ using an empirical estimator of the following expectation:
$Z({\bm{q}})=\operatorname{\mathbb{E}}_{\bm{\sigma}}\operatorname{\mathbb{E}}_{p_{v}({\bm{o}}|{\bm{q}};{\bm{\sigma}})}\Big[\frac{p_{\mathrm{ref}}({\bm{o}}|{\bm{q}};{\bm{\sigma}})}{p_{v}({\bm{o}}|{\bm{q}};{\bm{\sigma}})}\mathrm{e}^{r({\bm{q}},{\bm{o}})/\alpha}\Big].$
@@
Sequence log probability $\log p_{v}({\bm{o}}|{\bm{q}})$
Unlike the case of LLM, the exact sequence log probability is intractable in dLLM due to the presence of expectation over the random order $\bm{\sigma}$}. However, similar to the training of dLLM, we can leverage the negative ELBO 2 as a surrogate. Combined with the importance weight $w({\bm{o}}|{\bm{q}};{\bm{\sigma}})$, we introduce the weighted denoising cross-entropy (WDCE) loss for dLLM policy distribution matching:
$\min_{\theta}\operatorname{\mathbb{E}}_{{\bm{q}}\sim\mathcal{D}}\operatorname{\mathbb{E}}_{{\bm{\sigma}}}\operatorname{\mathbb{E}}_{p_{v}({\bm{o}}|{\bm{q}};{\bm{\sigma}})}\Big\{w({\bm{o}}|{\bm{q}};{\bm{\sigma}})\operatorname{\mathbb{E}}_{m\sim\operatorname{Unif}\{1,...,|{\bm{o}}|\}}\Big[\frac{|{\bm{o}}|}{m}\operatorname{\mathbb{E}}_{\mu_{m}({\widetilde{\bm{o}}}|{\bm{o}})}\sum_{d:\widetilde{o}_{d}=\mathsf{M}}-\log{\bm{\pi}}_{\theta}({\widetilde{\bm{o}}}|{\bm{q}})_{d,o_{d}}\Big]\Big\}.$ (11)
Notably, this loss highly resembles the DCE loss used in pre-training and the supervised fine-tuning (SFT) phase of dLLM. One major difference is that instead of using i.i.d. samples from $p_{*}$, we use importance sampling to weight samples from $p_{v}$ and obtain a valid training objective with theoretical guarantees. WDCE differs significantly from other popular RL training techniques such as PPO/GRPO in the following two key aspects:
@@
WDCE is an off-policy loss
The WDCE loss remains valid as the model parameter $\theta$ gets updated, since both the sampling policy $p_{v}$ and the important sampling target policy $p_{*}$ are independent of the current model policy $p_{\theta}$. This allows us to save generated rollouts in a replay buffer and reuse them for multiple training updates, without worrying excessively about numerical instability, leading to improved sample efficiency. On the other hand, for on-policy methods, to use a replay buffer, one would need to estimate importance weights with respect to the current model policy $p_{\theta}({\bm{o}}|{\bm{q}})$, i.e., $\frac{p_{\theta}({\bm{o}}|{\bm{q}})}{p_{v}({\bm{o}}|{\bm{q}})}$. Different from the case of LLM where such estimation can be done in one model forward pass, an accurate estimation in dLLM per training update is expensive, rendering the on-policy method less efficient. Moreover, for large models, where rollout generation and sequence likelihood estimation are typically handled by different implementations (such as vLLM and FSDP), this could lead to more nuanced, hard-to-detect biases that secretly undermine the algorithm’s performance (Yao et al., 2025). With WDCE, we are largely free of such concerns.
WDCE is a forward loss
Unlike the GRPO-style of algorithms that typically require keeping track of the entire rollout trajectories, WDCE leverages the forward (noising) process in training, which is a characteristic unique to diffusion LLMs. Once we obtain the final samples and their associated weights, we can discard the trajectories and perform training using the cheap forward process by randomly masking the data. This implies that the training speed when using WDCE largely depends on the model inference speed. With the advances of dLLM efficient inference techniques such as fast decoding algorithms and KV-cache techniques (Ma et al., 2025; Hu et al., 2025; Wu et al., 2025; Liu et al., 2025b), WDCE could also enjoy a great boost in efficiency. This method also effectively utilizes dLLM’s potential in surpassing LLMs in inference throughput, distinguishing it from other RL baselines that merely adapt LLM algorithms to dLLM. We defer a more detailed discussion of such properties to Sec. B.2.
Finally, we remark that while we developed WDCE through the lens of policy distribution matching learning 6, it can also be derived from the perspective of stochastic optimal control (SOC) as in Wang et al. (2025a); Zhu et al. (2025g), which we detail in Sec. B.1. In particular, we emphasize that the ELBO approximation of sequence-level log probabilities in matching KL divergence in 9 is equivalent to precisely matching path-level probabilities in the SOC framework, justifying the validity of such approximation.
3.3 Effective Training with Negative Gradient Insertion
While theoretically, minimizing the WDCE loss 11 probably leads to convergence of the model sequence distribution to $p_{*}({\bm{o}}|{\bm{q}})$, this could face practical issues due to the often limited number of rollouts generated per prompt. Ideally, we would want to promote the likelihood of “good” responses while decreasing those of “bad” responses. However, with WDCE, any response $\bm{o}$} will be associated with a positive weight $w({\bm{o}}|{\bm{q}};{\bm{\sigma}})$ due to the softmax operation, which may lead to ineffective learning in the low-batch-size scenario.